# Intention Analysis Makes LLMs A Good Jailbreak Defender

### v3

@@ 4.1 Setup
For safety datasets, we experiment on three main categories of jailbreak attacks, including three representative complex and stealthy in-the-wild jailbreak datasets (i.e. DAN (Shen et al., 2023), SAP200 (Deng et al., 2023a), and DeepInception (Li et al., 2023)), two popular optimization-based automatic jailbreak methods (i.e. GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023a)), and two advanced attacks for GPT-3.5 (i.e. multilingual attack called MultiJail (Deng et al., 2023b) and encryption-based attack named CipherChat (Yuan et al., 2024)). Besides, to evaluate the effect of our $\mathbb{IA}$ on helpfulness for general harmless in-distribution queries, we conduct experiments on three widely recognized datasets, i.e., AlpacaEval (Dubois et al., 2023), MMLU (Hendrycks et al., 2021) and TruthfulQA (Lin et al., 2022).
Figure 3: The MultiJail (under two scenarios) and CipherChat Datasets results on GPT-3.5 with and without our $\mathbb{IA}$. (a) Results on direct MultiJail dataset including Chinese (zh), Italic (it), Vietnamese (vi), Arabic (ar), Korean (ko), Thai (th), Bengali (bn), Swahili (sw), and Javanese (jv). (b) Results on malicious jailbreak prompt attached to MultiJail. (c) Results on CipherChat including ASCII (en), UTF (zh), Unicode (zh), and SelfCipher (en and zh) encryptions.
@@
For safety assessment, we annotate the harmfulness of responses and report attack success rate (ASR, Shen et al., 2023), where lower scores indicate stronger safety. Specifically, for DAN dataset, considering the complexity of responses, we adopt gpt-3.5-turbo-0613555https://openai.com/blog/chatgpt as the auto-annotation function following Deng et al. (2023a) and carry our human evaluation in Appendix C.1 to ensure the credibility. For the rest safety datasets, we annotate harmfulness following Zou et al. (2023) by matching refusal strings (e.g., “I’m sorry”; see Appendix C.2 for detailed settings).
@@
To evaluate the effectiveness of our $\mathbb{IA}$ method, we experiment on models varying scales and levels of alignment, including not only SFT models, i.e. Vicuna-7B/13B-v1.1 (Chiang et al., 2023) and MPT-30B-Chat (Team, 2023), but also RLHF models, i.e. ChatGLM-6B (Zeng et al., 2023), LLaMA2-7B-Chat (Touvron et al., 2023), and DeepSeek-67B-Chat (DeepSeek-AI, 2024). Beyond these representative open-source models, our experimentation extends to an advanced closed-source model, GPT-3.5 (gpt-3.5-turbo-1106) (OpenAI, 2023), which is renowned for its superior capabilities, including safety.
@@
We compare our $\mathbb{IA}$ with vanilla LLMs (without defense) and six popular defense methods, i.e., ICD (Wei et al., 2023b), (System-Mode) Self-Reminder (Xie et al., 2023), SmoothLLM (Robey et al., 2023), BPE-dropout (Jain et al., 2023), Self Defense (Helbling et al., 2023), and Moral Self-Correction (Ganguli et al., 2022). Among them, we report the first three representative defense methods in Table 1 and the entire in Table 9. For a fair comparison, we closely follow the best default parameters in their papers.
@@
The detailed $\mathbb{IA}$ prompts for experiments are provided in Figure 8666To assess the resilience of our method against specific expressions, we construct alternative $\mathbb{IA}$ prompts and conduct comparison experiment in Section 5.2.. For the DAN dataset, we compile an evaluation dataset of 1560 samples by extracting 195 questions from each jailbreak community within the forbidden question set (Shen et al., 2023). For GCG, we follow Zou et al. (2023) and conduct transfer attacks on Vicuna-7B and 13B. The adversarial suffix achieving the lowest loss after 500 steps of optimization are adopted to further attack target models on 100 individual harmful behaviors (Wei et al., 2023b). For open-source models, we download them from HuggingFace [^7] https://huggingface.co/models. For closed-source models, we obtain the responses of GPT-3.5 via API calls. Throughout our experiments, we set a temperature of zero for deterministic outcomes Peng et al. (2023) and a generation length of 1024 tokens, employing default system prompt templates for each LLM regarding their official reports.
4.2 Main Results
Our method significantly boosts safety across a diverse range of LLMs and consistently outperforms other defense baselines.
In Table 1, we represent the ASR of varying defense methods on different LLMs under various jailbreak attacks. We can observe that: 1) $\mathbb{IA}$ effectively reduces ASRs across a diverse range of LLMs, irrespective of their inherent safety levels. For LLMs with high vanilla ASRs, such as ChatGLM-6B, Vicuna-7B, Vicuna-13B, MPT-30B-Chat, and DeepSeek-67B-Chat, we significantly lower the average ASRs from 72.7% to 3.79%. Similarly, for LLMs presenting lower vanilla ASRs, such as LLaMA2-7B-Chat and GPT-3.5, $\mathbb{IA}$ further reduces their average ASRs from 13.73% to mere 0.1%. 2) $\mathbb{IA}$ consistently surpasses all defense baselines. For example, on Vicuna-7B, Vicuna-13B, and MPT-30B-Chat, $\mathbb{IA}$ significantly averages a reduction of 37.4% in ASR compared to the second best baseline. We conjecture such satisfactory performance to be attributed to our design of intention analysis mechanism, which leverages LLM’s awareness of harmful content to enhance safety against varying jailbreak attacks.
@@ 4.2 Main Results
Ours 73.5 48.3 38.2 55.1
Table 2: General performance on helpful dataset upon different models in terms of Win Rate (%) for AlpacaEval, Accuracy (%) for MMLU and MC1, MC2 (%) for TruthfulQA.
Recent studies Deng et al. (2023b); Yong et al. (2023) reveal that the multilingual jailbreak poses a new defense challenge for LLMs. Yuan et al. (2024) and Wei et al. (2023a) also emphasize the struggles of more powerful LLMs, such as GPT-3.5, to stay safe when countering encryption-based attack. To verify the effectiveness of our method in these advanced jailbreak scenarios, we reproduce MultiJail and CipherChat following Deng et al. (2023b) and Yuan et al. (2024), respectively, and conduct further experiments on GPT-3.5888We observe a high rate of invalid responses on smaller LLMs like ChatGLM-6B, Vicuna-7B and Vicuna-13B under MultiJail, as demonstrated in Deng et al. (2023b). And smaller LLMs may lack advanced encryption capability, which is required by CipherChat (Yuan et al., 2024). So we only adopt GPT-3.5 for these advanded jailbreak attacks experiments.. The results of GPT-3.5 with and without our $\mathbb{IA}$ are presented in Figure 3. We observe that 1) our $\mathbb{IA}$ consistently maintains performance in low-resource languages, e.g., th, bn, sw, and jv, even in scenarios where a malicious jailbreak prompt [^9] We adopt the same jailbreak prompt in Deng et al. (2023b), namely AIM. is attached to the multilingual attacks, 2) our $\mathbb{IA}$ significantly enhances safety when facing advanced encryption-based attack, even under the most effective SelfCipher attack. These demonstrate the effectiveness of our intention analysis defense mechanism under more advanced jailbreak attacks.
Our method maintains LLM’s general helpfulness on harmless queries.
To explore the impact of our method on the general performance of LLMs, we conduct experiments on several acknowledged helpfulness datasets and report the results in Table 2. As observed, for harmless user prompts, our $\mathbb{IA}$ does not compromise the general helpfulness, indicated by a comparable win rate on AlpacaEval, a similar accuracy on MMLU, and better MC1 and MC2 scores on TruthfulQA compared with vanilla LLMs. These results indicate that $\mathbb{IA}$ can be deployed in real applications to enhance LLM safety while preserving general helpfulness.
Our method ensures the helpfulness of safe responses for harmful queries.
To comprehensively study the impact of our $\mathbb{IA}$ on responses to harmful queries, we conduct a manual review of 100 random-sampled refusals on both DAN and SAP200 datasets for each of the seven LLMs under our $\mathbb{IA}$. We manually check for 1) reasonable explanations about why it rejects to respond and 2) positive suggestions for the user. In Table 3, our findings show that, following $\mathbb{IA}$, almost all LLMs can give detailed explanations in nearly 100% of cases (except for GPT-3.5 due to OpenAI’s external moderation mechanism during API interactions (OpenAI, 2023)). Specifically, we find LLMs indeed conduct more nuanced handling regarding varying unsafe intents (e.g. terrorism and violence). Detailed cases are presented in Figures 14.
LLaMA2-7B-Chat Vicuna-7B Vicuna-13B
@@
SAP200 100% 100% 49%
5 Discussion
5.1 Analysis of Intention Analysis Mechanism
Most LLMs can effectively recognize the intention behind jailbreak queries.
@@ 5.1 Analysis of Intention Analysis Mechanism
Figure 5: Performance of $\mathbb{IA}$ with varying correct intention ratio on DAN dataset.
@@
We observe that most LLMs are highly effective in recognizing intentions behind complex and stealthy jailbreak queries, achieving a nearly 100% success rate in Vicuna-13B, MPT-30B-Chat, and DeepSeek-67B-Chat. Particularly, the intention recognition rate of LLaMA2-7B-Chat is relatively lower due to its excessively strong inherent safety, leading to directly refusing to answer the user query (see Figure 15 for detailed cases).
Two factors that affect $\mathbb{IA}$ performance are intention analysis outcomes and inherent LLM safety.
(1) Intention analysis outcomes [^11] We also conduct cross-intention analysis experiment on Vicuna-7B and Vicuna-13B in Appendix F to explore the effect of different intention analysis LLMs.: In Figure 5, we explore the influence of intention analysis outcomes by simulating varying correct intention ratios through replacing the gold intention identified in the first stage response with an invalid field (such as “[secret]”). We can see that $\mathbb{IA}$ performance gradually improves as the correct intention ratios increase, indicating that enhancing the intention analysis ability can further boost $\mathbb{IA}$ effectiveness [^12] Despite a 0% correct intention ratio, $\mathbb{IA}$ remains effective compared to vanilla LLMs, since the intention analysis sequence format allows replacing gold intention labels with invalid ones to be marginally detrimental, as recognized by the In Context Learning (ICL) community (Min et al., 2022)..
Methods Safety Helpfulness
SAP200 DAN AlpacaEval
Vicuna-7B 73.4 44.3 66.2
[…]
Vicuna-7B + Ours 0.31 2.89 62.8
(2) Inherent LLM safety: In Figure 4, among LLMs achieving nearly 100% in intention recognition rates, there is a notable difference in harmful response rates, ranging from 0.3% for Vicuna-7B to 19.3% for MPT-30B-Chat, indicating the impacts of inherent LLM safety on $\mathbb{IA}$ results (see Figure 16 for related case study). These pose two potential directions for further improving $\mathbb{IA}$ performance in the future: enhancing the LLMs’ intention analysis capability and their inherent safety.
5.2 Efficiency and Robustness of Our $\mathbb{IA}$
Our method achieves comparable safety with well safety-trained LLMs without need for additional training.
Our method aims to enhance LLM safety in the inference stage. A natural question arises: how does its performance compare to well safety-trained LLMs? To answer this, we compare our method with a representative well safety-trained LLM, i.e., LLaMA2-7B-Chat. The results are listed in Table 4. We can see that our method achieves comparable performance to LLaMA2-7B-Chat on safety datasets while outperforming LLaMA2-7B-Chat on the helpfulness dataset by almost 5%. This demonstrates the advantage of our $\mathbb{IA}$ to achieve both safety and helpfulness goals without additionally resource-consuming safety training.
Our efficient one-pass variant of $\mathbb{IA}$ still works well.
@@ 5.2 Efficiency and Robustness of Our $\mathbb{IA}$
As aforementioned, to maximize the performance, our $\mathbb{IA}$ follows a two-stage process. A natural question arises of whether our mechanism can be merged into one step, to save the decoding overhead. To verify this, we design a cheaper one-pass $\mathbb{IA}$, without letting the LLMs separately explain the intention (see Figure 9 for detailed prompts). As shown in Table 5, we see that 1) both the two-stage and one-pass $\mathbb{IA}$ perform well on different LLMs, 2) two-stage $\mathbb{IA}$ can further improve LLM safety compared with one-pass $\mathbb{IA}$.
Our method is robust to different $\mathbb{IA}$ prompts.
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions of our $\mathbb{IA}$ (see Figure 10 for detailed prompts) and conduct comparison with the prompt set we used in main experiments. Table 6 presents the comparison results. We observe that three prompt sets all result in significant and comparable decreases in ASRs on the SAP200 across different LLMs. This uniformity highlights a key conclusion: it is the intention analysis mechanism, but not the specific instruction expressions, that makes our $\mathbb{IA}$ effective.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 73.4 65.4 45.8
+ One-Pass $\mathbb{IA}$ 5.50 1.13 39.0
+ Two-Stage $\mathbb{IA}$ 0.31 1.12 6.12
@@
Table 5: Comparison of our $\mathbb{IA}$ with different implementations (one-pass and two-stage) on SAP200 in terms of ASR (%). The best are highlighted in bold.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 73.4 65.4 45.8
+ Prompt A 2.94 0.88 5.81
+ Prompt B 5.13 2.06 5.44
+ Ours 0.31 1.12 6.12

### v4

@@ 4.1 Setup
For safety datasets, we experiment on three main categories of jailbreak attacks, including three representative complex and stealthy in-the-wild jailbreak datasets (i.e. DAN (Shen et al., 2023), SAP200 (Deng et al., 2023a), and DeepInception (Li et al., 2023)), two popular optimization-based automatic jailbreak methods (i.e. GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023a)), and two advanced attacks for GPT-3.5 (i.e. multilingual attack called MultiJail (Deng et al., 2023b) and encryption-based attack named CipherChat (Yuan et al., 2024)).
Besides, to evaluate $\mathbb{IA}$’s effect on helpfulness for general benign queries, we conduct experiments on three widely recognized datasets, i.e., AlpacaEval (Dubois et al., 2023), MMLU (Hendrycks et al., 2021) and TruthfulQA (Lin et al., 2022).
@@
For safety assessment, we annotate the harmfulness of responses and report attack success rate (ASR, Shen et al., 2023), where lower scores indicate stronger safety. Specifically, for DAN dataset, considering the complexity of responses, we adopt gpt-3.5-turbo-0613666https://openai.com/blog/chatgpt as the auto-annotation function following Deng et al. (2023a) and carry our human evaluation in Appendix C.1 to ensure the credibility. For other safety datasets, we annotate harmfulness following Zou et al. (2023) by matching refusal strings (e.g., “I’m sorry”; see Appendix C.2 for detailed settings).
@@
To evaluate $\mathbb{IA}$’s effectiveness, we experiment on representative LLMs with varying scales and alignment levels, including not only SFT models, i.e. Vicuna-7B/13B-v1.1 (Chiang et al., 2023) and MPT-30B-Chat (Team, 2023), but also RLHF models, i.e. ChatGLM-6B (Zeng et al., 2023), Llama2-7B-Chat (Touvron et al., 2023), Llama3-8B-Instruct [^7] https://ai.meta.com/blog/meta-Llama-3/, and DeepSeek-67B-Chat (DeepSeek-AI, 2024). Beyond open-source LLMs, our experimentation extends to an advanced closed-source LLM, GPT-3.5 (gpt-3.5-turbo-1106) (OpenAI, 2023), renowned for its superior capabilities, especially safety alignment.
@@
We compare our $\mathbb{IA}$ with vanilla LLMs (no defense) and seven popular defense methods, i.e., Input Check [^8] We create an Input Check baseline by using the prompt in Helbling et al. (2023) and operate in the input space to let LLMs judge whether a query is harmful or not., ICD (Wei et al., 2023b), (System-Mode) Self-Reminder (Xie et al., 2023), SmoothLLM (Robey et al., 2023), BPE-dropout (Jain et al., 2023), Self Defense (Helbling et al., 2023), and Moral Self-Correction (Ganguli et al., 2022). The first four representative defense methods are reported in Table 1 and others in Table 7 in Appendix due to page limitation. Besides, a training method is also included in Appendix E.3 and results show $\mathbb{IA}$ achieves both safety and helpfulness goals without additional resource-consuming safety training. For a fair comparison, we closely follow the best default parameters in their papers.
@@
The detailed $\mathbb{IA}$ prompts for experiments are provided in Figure $9$To assess the resilience of our method against specific expressions, we construct two alternative $\mathbb{IA}$ prompts in Appendix D.2 and experiment results demonstrate $\mathbb{IA}$’s effectiveness is irrespective of specific expressions.. For the DAN dataset, we compile an evaluation dataset of 1560 samples by extracting 195 questions from each jailbreak community within the forbidden question set (Shen et al., 2023). For GCG, we follow Zou et al. (2023) and conduct transfer attacks on Vicuna-7B and 13B. The adversarial suffix achieving the lowest loss after 500 steps of optimization are adopted to further attack target models on 100 individual harmful behaviors (Wei et al., 2023b). For open-source models, we download them from HuggingFace [^10] https://huggingface.co/models. For closed-source models, we obtain the responses of GPT-3.5 via API calls. Throughout our experiments, we set a temperature of zero for deterministic outcomes Peng et al. (2023) and a generation length of 1024 tokens, employing default system prompt templates for each LLM regarding their official reports. All experiments are carried out on a solitary node outfitted with 8 A100-SXM80GB GPUs.
@@
[…]
Table 2: General performance on helpful dataset upon different models in terms of Win Rate (%) for AlpacaEval, Accuracy (%) for MMLU and MC1, MC2 (%) for TruthfulQA. “—” means lacking official implementation for distributed larger models or white-box LLM weights required.
4.2 Main Results
Performance of safety on various jailbreak attacks
In Table 1, we represent the ASR of several defense baselines on different LLMs under various jailbreak attacks as well as inference time comparison [^11] Due to memory constraints, the Deepspeed Zero-3 algorithm was employed for larger models, MPT-30B and DeepSeek-67B, resulting in relatively longer inference times.. We can observe that: 1) $\mathbb{IA}$ effectively reduces ASRs across a diverse range of LLMs along with an acceptable time cost. For LLMs with high vanilla ASRs, such as ChatGLM-6B, Vicuna-7B, Vicuna-13B, MPT-30B-Chat, and DeepSeek-67B-Chat, we significantly lower the average ASRs from 72.7% to 3.79%. Similarly, for LLMs presenting lower vanilla ASRs, such as Llama2-7B-Chat, Llama3-8B-Instruct, and GPT-3.5, $\mathbb{IA}$ further reduces their average ASRs from 13.8% to mere 0.1%. 2) $\mathbb{IA}$ maintain its effectiveness even in scenarios where other defense methods struggle. For example, AutoDAN leverages LLMs to automatically attack based on optimization and thus is hard to defend. While the baselines have ASRs of at least 83% on Vicuna-7B and 13B under AutoDAN, $\mathbb{IA}$ can still provide significant defense with a low ASR of under 11%. Notably, $\mathbb{IA}$ can also integrate with another defensive method to enhance performance but with additional computation overhead (see Appendix E.4 for details). Moreover, we also extend to more advanced jailbreak attacks including multilingual and encryption-based attacks, and demonstrate our consistent effectiveness on ChatGPT (see Appendix E.1). Further analysis regarding our good performance will be discussed in Section 5.
Performance of general helpfulness for benign queries
An effective defense method is expected to maintain general abilities as well. To explore the impact of our method on the general performance of LLMs, we conduct experiments on several acknowledged helpfulness datasets and report the results in Table 2. As observed, for harmless user prompts, our $\mathbb{IA}$ does not significantly compromise the general helpfulness on AlpacaEval, MMLU, and TruthfulQA benchmarks compared with vanilla LLMs. These results indicate that $\mathbb{IA}$ can be deployed in real applications to enhance LLM safety while preserving general helpfulness. More comparison results with other defensive methods can be found in Table 11 in Appendix. To further study $\mathbb{IA}$’s impact on LLM’s helpfulness, we also conduct both manual and automatic checks about safe refusal’s helpfulness for harmful queries and find that $\mathbb{IA}$ enables LLMs to effectively give safe refusals with satisfactory helpfulness for harmful queries, instead of simple rejection (see Appendix G for detailed analysis).
5 Discussion of $\mathbb{IA}$ Mechanism
5.1 Can LLMs successfully generate the intentions behind jailbreak queries?
@@
We observe that: 1) Most LLMs are highly effective in recognizing intentions behind complex and stealthy jailbreak queries, achieving a nearly 100% success rate in Vicuna-13B, MPT-30B-Chat, and DeepSeek-67B-Chat. Particularly, the intention recognition rate of Llama2-7B-Chat is relatively lower due to its excessively strong inherent safety leading to direct refusals to harmful user queries [^13] We do not present Llama3-8B-Instruct for the same reason that its strong inherent safety leads to almost all direct refusals in the intention analysis stage.(see Figure 17 for detailed cases). 2) In adversarial scenarios, it is easier for most LLMs to generate intentions than directly generate safe responses. Setting the SAP dataset as an example, most LLMs can successfully identify more than 90% of the adversarial intents. While in Table 1, they can only generate averagely $\sim$30safe responses.
Figure 4: Performance of $\mathbb{IA}$ with varying correct intention ratio on DAN dataset. From left to right: the correct intentions are replaced with masked and random intention, respectively.
5.2 What if LLMs generate incorrect intentions?
To understand the effect of intention analysis errors, we examine two extreme cases: 1) recognized intentions are masked with an invalid field (e.g., “[secret]”); 2) recognized intentions are replaced with randomly sampled tokens from LLM’s vocabulary, simulating a severely wrong case. Figure 4 shows $\mathbb{IA}$’s performance across different correct intention ratios on DAN dataset. Overall, $\mathbb{IA}$’s performance declines with increasing intention errors but consistently maintains a much lower ASR (below 10%) compared to the vanilla baseline, indicating $\mathbb{IA}$’s some robustness to wrong intentions.
Notably, $\mathbb{IA}$ remains effective even at a 0% correct intention ratio. This can be attributed to the role of the intention analysis sequence format, allowing replacing true intentions with invalid ones to be marginally detrimental, as widely recognized by the In Context Learning (ICL) community (Min et al., 2022). Further explanation can be found at Appendix F. However, exploring the underlying principles of how sequence formats affect outcomes is beyond the scope of this work.
5.3 What is the underlying principle of $\mathbb{IA}$?
This section explores how $\mathbb{IA}$ works by analyzing the model’s attention distribution across different prompt components during response generation [^14] Inspired by Wang et al. (2023), the attention score is calculated by averaging the maximum attention scores for each prompt component across the samples in the DAN dataset. (see Figure 5) [^15] As depicted in Figure 2, vanilla prompt consists of jailbreak prompt and harmful question. IA-Stage 1 prompt adds an intention analysis instruction before the vanilla prompt. IA-Stage 2 then combines the IA-Stage 1 prompt, the recognized intention from Stage 1, and the final response instruction.. As shown, the model under vanilla prompt pays significant attention to the jailbreak prompt, leading to potentially harmful responses. In contrast, $\mathbb{IA}$ at both stages significantly reduces LLM’s attention to the jailbreak prompt while increasing attention to user intent, making LLM less likely to follow jailbreak prompts and leading to safer responses.
To further illustrate $\mathbb{IA}$’s effect, Figure 6 presents a layer-wise comparison of attention on the jailbreak prompt between the vanilla and $\mathbb{IA}$ prompts. The results show that $\mathbb{IA}$ consistently reduces the model’s attention on the jailbreak prompt across all layers, further indicating $\mathbb{IA}$’s effectiveness in suppressing LLM’s tendency to follow jailbreak prompts.
Figure 5: Comparison of Vicuna-13B’s attention scores on different prompt components of different methods. The average attention score is computed on DAN dataset. $\mathbb{IA}$ largely decreases model’s attention to jailbreak prompt (red bar) in both two stages.
6 Further Discussion
Two factors influence $\mathbb{IA}$ performance. (1) Intention analysis ability: As shown by the solid lines in Figure 4, $\mathbb{IA}$ performance improves with higher correct intention ratios, suggesting that better intention analysis ability can further enhance effectiveness [^16] We also conduct cross-intention analysis experiment on Vicuna-7B and Vicuna-13B in Appendix H to explore the effect of different intention analysis LLMs.. (2) Inherent LLM safety: Figure 3 shows that even among LLMs with nearly 100% intention recognition rates, the final harmful response rates vary notably—from 0.3% for Vicuna-7B to 19.3% for MPT-30B-Chat—highlighting the impact of inherent LLM safety on $\mathbb{IA}$ results (see Figure 18 for a related case study). These suggest two improvement directions: enhancing LLMs’ intention analysis ability and their inherent safety.
Our efficient one-pass variant of $\mathbb{IA}$ provides a more cost-effective choice.
@@ 6 Further Discussion
As aforementioned, to maximize the performance, our $\mathbb{IA}$ follows a two-stage process. A natural question arises of whether our mechanism can be merged into one step, to save the decoding overhead. To verify this, we design a cheaper one-pass $\mathbb{IA}$ variant (see Figure 10 for detailed prompts). From results in Table 3, we see that: 1) For more powerful models, such as Vicuna-7B and 13B, one-pass $\mathbb{IA}$ achieves comparable performance to two-stage $\mathbb{IA}$ in a more cost-effective manner. 2) For less powerful models, i.e., ChatGLM-6B, one-pass $\mathbb{IA}$’s effectiveness diminishes to some extent. In such cases, two-stage $\mathbb{IA}$ is necessary to sustain satisfactory performance.
Figure 6: Comparison of Vicuna-13B’s attention scores on jailbreak prompt between Vanilla and $\mathbb{IA}$ methods across different model layers. The average attention score is computed on DAN dataset. High scores means greater influence of jailbreak prompt on the generated response.
Vicuna-7B Vicuna-13B ChatGLM-6B Time Cost
Vanilla 73.4 65.4 45.8 13.2
+ One-Pass $\mathbb{IA}$ 5.50 1.13 39.0 12.6
+ Two-Stage $\mathbb{IA}$ 0.31 1.12 6.12 19.5
@@
Table 3: Comparison of our $\mathbb{IA}$ with different implementations (one-pass and two-stage) on SAP200 in terms of ASR (%) and average Time Cost (s/sample).